OpenStudy (safyousuff97):

Given the function f(x) = 2(3)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (safyousuff97):

@rvc

rvc (rvc):

@mathmate

OpenStudy (mathmate):

I suspect the function is meant to be \(2(3^x)\) in which case the rates of change would be different between the two sections.

OpenStudy (safyousuff97):

yes except 2(3) x on the top

OpenStudy (mathmate):

average rate of change is given by \(\Delta y/\Delta x\), so for the first section, average rate of change =\((f(1)-f(0))/(1-0)=(2*3^1-2*3^0)/(1-0)=4\) You can calculate the average rate of change for the other interval similarly

OpenStudy (safyousuff97):

i'm sorry i'm confused can u explain what u just did?

OpenStudy (safyousuff97):

@mathmate

OpenStudy (caozeyuan):

top is f(end)-f(start), bottom is end-start

Nnesha (nnesha):

what a coincidence.. what grade are u in??

OpenStudy (safyousuff97):

9th

Nnesha (nnesha):

cool.

Nnesha (nnesha):

for section A first you need to find y (or find f(1) ,f(0)values when x=1 ,x=0 substitute x for 1 and 0 solve for y \[\rm f(1)=2(3)^1 \] what would you get when u substitute x for 1 ??? f(1)= ?